OpenStudy (anonymous):

help plzz Name the Quadratic, Linear, and Constant Terms of f(x)=2(5x-3)

OpenStudy (anonymous):

plzz answer its algebra 2

OpenStudy (anonymous):

ok multiply the 2 inside \[\Large f(x)=10x-6\]

OpenStudy (anonymous):

is there any quadratic term?

OpenStudy (anonymous):

@pitachips check your question ? i think it should be 2x(5x-3) is it?

OpenStudy (anonymous):

nope

OpenStudy (anonymous):

ok then there is no Quadratic term linear term is 10x and constant -6

OpenStudy (anonymous):

@pitachips did you get it?

OpenStudy (anonymous):

yes thanxks but i have another question 1. f(x )=(2x+1)(5x-7): classify if its quadratic, linear, or neither and 2.simplify f(x)=-4(2x^2+3x-3)

OpenStudy (anonymous):

Both are Quadratic i already told you if there is any \[\Large x^2\] term in the equation then it is Quadratic. in the 2) it is obviously Quadratic because it has x^2 term for first one first Multiply them then simplify .you will get x^2 term i hope it is clear!

OpenStudy (anonymous):

but for number 2. f(x )=-4(2x^2+3x-3)= -8x^2-12x-12????

OpenStudy (anonymous):

yes you are right:) but slight mistake (after multiplication it should be +12) f(x )=-4(2x^2+3x-3)= -8x^2-12x+12

OpenStudy (anonymous):

THX